] Use a proportion to solve the problem.
The ratio of the distances a 7-iron and a 5-iron will drive a golf ball is 5 to 6. If a golfer averages 137 yards with a 7-iron, how far should he average with a 5-iron?

Answer by checkley77(12844)

5/6=x/137
6x=5*137
6x=685
x=685/5
x=114.17 yards.